OpenStudy (anonymous):

Find the slope-intercept form of the equation of a line with slope -1/4 and y-intercept (0,-2).

OpenStudy (anonymous):

Can Someone give me the answer for this problem?

OpenStudy (asnaseer):

do you know what the slope-intercept form is?

OpenStudy (anonymous):

Find the slope-intercept form of the equation of a line with slope -1/4 and y-intercept (0,-2). Question 1 options: y= -1/4x-2 y= -2x- 1/4 y= -1/4x+2 y= -1/4x

OpenStudy (anonymous):

Those are the choices

OpenStudy (asnaseer):

I was asking if you knew what slope-intercept form means?

OpenStudy (anonymous):

yes i do

OpenStudy (asnaseer):

so what is the general form of a line in slope-intercept form?

OpenStudy (anonymous):

i already did this problem im just trying to get an answer from someone else to see if its the same answer as me

OpenStudy (asnaseer):

what answer did you get?

OpenStudy (anonymous):

y= -1/4x-2

OpenStudy (asnaseer):

that is correct

OpenStudy (asnaseer):

you should probably phrase questions like these as "I get this answer ... - can please someone check if this is correct?"